• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Jack

    Log entry time 17:48:56 on June26,2000

    Entry number 44921

    keyword=Gas flow controllers replaced.

    The argon mass flow controller seems to have been out of tolerance. Both
    argon and ethane controllers have been replaced with units of a more recent
    calibration. They seem to be functioning correctly so far. The set and
    readback voltages in EPICS are being logged. While the moller runs we will monitor
    the controllers' performance.